# Finite-temperature Rydberg arrays: quantum phases and entanglement characterization

### v1

@@ Abstract
With one or more electrons in a highly excited energy state, Rydberg atoms display strong interparticle interactions Urban et al. (2009); Löw et al. (2012); Weber et al. (2017), making them excellent candidates for programmable quantum devices Wu et al. (2021); Browaeys and Lahaye (2020). With the possibility of arranging individual Rydberg atoms in arrays of arbitrary geometries using optical tweezers Kaufman and Ni (2021); Barredo et al. (2018); Endres et al. (2016); Manetsch et al. (2024), recent advancements in quantum computing with neutral atoms have demonstrated high-fidelity qubit gates Levine et al. (2019); Zhuo et al. (2022); Evered et al. (2023) and arbitrary connectivity Bluvstein et al. (2022, 2023). Moreover, the progress in analog quantum simulation enables the preparation and control of hundreds of neutral atom qubits, allowing to probe quantum many-body properties directly through measurements Gross and Bloch (2017); Browaeys and Lahaye (2020); Wurtz et al. (2023). Quantum simulators have demonstrated to be successful in exploring the various quantum phases, critical behavior, and underlying universal properties Bernien et al. (2017); Keesling et al. (2019); De Léséleuc et al. (2019); Scholl et al. (2021); Ebadi et al. (2021); Zhang et al. (2024).
@@
The paper is structured as follows: in Section I, we explain the numerical method. Then in Section II, we introduce the Rydberg Hamiltonian that we simulate, and give a short overview of the previous findings regarding the structure of the ground state phase diagram. We show our results in Sections III and IV. Finally, we summarize the conclusions in Section V. Details on the numerical method can be found in Appendices A and B.
@@ I Method
In order to construct an efficient numerical algorithm for studying finite-temperature physics, it is important to choose an appropriate representation of the density matrix. Since the TTO representation is well-suited for the extraction of the otherwise hardly accessible mixed-state properties in a low-temperature regime, here we develop the algorithm for constructing the TTO density matrix for quantum many-body systems at thermal equilibrium. In this section, we first provide a recap of TTOs and their advantages in representing and compressing the density matrix. Then, we present the main steps of our numerical algorithm.
@@ I.1 Tree Tensor Operator
where |$\ket{\psi_{j}}$ is the $j^{\mathrm{th}}$ eigenstate of the system’s Hamiltonian with corresponding eigenenergy $E_{j}$, $p_{j}$ are the classical probabilities of finding a system in a state |$\ket{\psi_{j}}$, $Z$ is the partition function, and $\beta=1/T$ with $T$ being temperature (in Boltzmann’s constant units, i.e. $k_{B}=1$). Even though, in principle, the number of eigenstates that contribute to the density matrix grows exponentially with the system size, the probabilities $p_{j}$ decay exponentially with the eigenstate’s energy. The temperature factor in the probabilities regulates the flatness of the distribution and one can see that only the low-energy states have a non-negligible contribution to the density matrix in the low-temperature regime (Fig. 1c). In the TTO structure, the middle connecting link corresponds to the sum over eigenstates in Eq. (1), and thus we have direct access to the number of states kept in the density matrix, determined by the corresponding bond dimension $K_{0}$ (Fig. 1a). With the ability to keep only the first $K_{0}$ eigenstates and discard the information about the rest, we keep the computational cost of TTO algorithms low, while retaining the physically accurate approximation. Altogether, we employ two separate bond dimension bounds in the TTO: $m$ as the bond dimension bounding system-to-system correlations and $K_{0}$ as the bound for the mixing bond dimension, i.e. for system-environment correlations.
@@
Figure 2: An algorithm for obtaining the thermal density matrix in the TTO form. An infinite-temperature density matrix is always proportional to an identity matrix, and it is therefore straightforward to write in the LPTN form. We take this infinite-temperature density matrix as an initial state and evolve it via the imaginary time evolution, i.e. consecutively applying the operator $e^{-\mathrm{d}\beta\hat{H}/2}$ on both sides of the LPTN. Here, $\mathrm{d}\beta/2$ is an imaginary time step which, in terms of physical interpretation, represents a small decrease in temperature. This way, we can reach a density matrix at an arbitrarily low temperature. Once we reach the desired temperature, we convert the LPTN into the TTO form with the procedure described in Appendix A.
I.2 Obtaining the Tree Tensor Operator density matrix
To access the system’s properties at an arbitrary temperature, we evolve the infinite-temperature density matrix according to the imaginary time evolution Verstraete et al. (2004); Zwolak and Vidal (2004). This is possible due to the fact that an infinite-temperature density matrix, $\hat{\rho}_{\infty}$, is proportional to an identity matrix, allowing to us write:
$\begin{split}\hat{\rho}&=e^{-\beta\hat{H}}=e^{-\beta\hat{H}/2}\mathds{1}e^{-\beta\hat{H}/2}\\ &\propto e^{-\beta\hat{H}/2}\hat{\rho}_{\infty}e^{-\beta\hat{H}/2}.\end{split}$ (2)
Therefore, obtaining the density matrix at a certain temperature is mathematically equivalent to performing a time evolution of the identity matrix with an imaginary timestep.
@@ I.2 Obtaining the Tree Tensor Operator density matrix
We evolve LPTN in imaginary time using the time-dependent variational principle (TDVP) Haegeman et al. (2011, 2016) algorithm. Once we obtain the finite-temperature LPTN density matrix, we convert it to a TTO with an iterative procedure described in Appendix A. The whole procedure for obtaining a TTO thermal density matrix is summarized in Fig. 2.
@@ II Rydberg atom Hamiltonian
To identify the ground state phases in the parameter range of our simulations, we measure the entanglement negativity $E_{N}$ between two halves of the system at temperature $T=0.025\ \Omega$ in Fig. 3. As shown in the analysis in Section III, this temperature is low enough for the system to be approximately in the ground state. The entanglement negativity is a mixed-state entanglement measure defined as
$\begin{split}&E_{N}=\frac{\lVert\hat{\rho}^{T_{N/2}}\rVert_{1}-1}{2},\end{split}$ (4)
where $\hat{\rho}^{T_{N/2}}$ is the partial transpose of $\hat{\rho}$ with respect to half of the system and $\lVert A\rVert_{1}$ is the trace norm, $\lVert A\rVert_{1}=Tr(\sqrt{A^{{\dagger}}A})$. Even though entanglement negativity is constructed as a mixed-state entanglement measure, here we show that it can also be used for quantifying the pure state entanglement, as the phase diagram in Fig. 3 is in line with the existing results obtained via the entanglement entropy Rader and Lauchli (2019); Yu et al. (2022); Maceira et al. (2022). The phase transition lines correspond to maxima in Fig. 3 where one can differentiate three phases in our parameter range: a disordered phase and two crystalline phases. It is known that the transition line between the $\mathds{Z}_{2}$ and disordered phase exhibits the Ising universality behavior. The transition between the $\mathds{Z}_{3}$ and the disordered phase is more complex, and in the parameter range of our simulations, it contains the non-conformal chiral transition line and an isolated point of the three-state Potts universality class.
@@ III Finite temperature phase diagrams
Recall that the accuracy of our method is limited by the number of excited states we keep in a density matrix, $K_{0}$. This means that, when the state becomes more mixed, the state truncation error becomes larger at fixed $K_{0}$. An easy way of estimating the error induced by discarding the excited states is by tracking the norm loss during the LPTN to TTO conversion. The blue lines in Fig. 4 indicate the temperatures after which the preserved norm falls under a certain factor. As expected, the error increases with the system size for a fixed temperature value. This is because the number of excited states in the spectrum grows exponentially with the system size, and thus, increasing the system size, the number of states $K_{0}$ will represent a smaller ratio of all possible states.
To get an overview of the temperature resilience of ground state phases over the entire phase diagram, we plot the threshold temperature at which the purity falls below a chosen cutoff value, across the entire Hamiltonian parameter range. The obtained finite-temperature phase diagrams for two different values of threshold purity value, $\gamma_{cut}=\{0.5,0.9\}$, are shown in Fig. 5 and one can see that they outline the features of a ground state phase diagram from Fig. 3. The plot minima correspond to the borders between the phases, and the ground state becomes more stable as we go deeper into the crystalline phases. Again, we observe that the transition from the ground state to a mixture of excited states occurs faster for larger systems, i.e. as we go towards the thermodynamic limit.
@@ IV Mixed-state entanglement at quantum critical points
In this section, we focus on the conformal phase transition points. In particular, we compute the mixed-state entanglement monotones at two points belonging to the Ising universality class and the Potts universality class Yu et al. (2022) (marked with a blue diamond and a green star in Fig. 3). Exploiting the TTO ansatz, we measure the temperature dependence of the bipartite entanglement between the two halves of the system via two mixed-state entanglement measures: the entanglement negativity defined in Eq. (4), and the entanglement of formation. The entanglement of formation is the mixed-state generalization of entanglement entropy. Mathematically, it is defined as a weighted sum of entanglement entropies minimized over all the possible pure-state decompositions of $\hat{\rho}$, and is defined as:
with $\mathcal{S}_{E}$ denoting the entanglement entropy of half-system bipartition. In general, calculating the entanglement of formation is an extremely challenging task, and exact solutions are known only for small system sizes Hill and Wootters (1997); Wootters (1998) or very specific cases Terhal and Vollbrecht (2000); Vollbrecht and Werner (2001).
Here, we compute it using the procedure presented in Arceci et al. (2021), improved as described in Appendix B. In order to improve the convergence of the results, we truncate the $K_{0}$ dimension even further and keep the first $K_{0}=10$ states in the density matrix for the computation of entanglement of formation. Note that, as discussed in Section III, reducing the number of states in the density matrix increases the error in the high-temperature region. Therefore, we carry out the entanglement of formation analysis only in the low-temperature region in which the norm loss is smaller than $1\%$.
@@
The next step is to analyze and quantify finite-size effects, i.e. how entanglement scales with system size and temperature. The numerical results on the entanglement of formation $E_{F}$ in low-temperature Ising and Luttinger liquid XXZ model Arceci et al. (2021) predicted that in the vicinity of the conformal quantum critical point, $E_{F}$ should follow the logarithmic scaling. This scaling relation is the finite-temperature extension of the conformal field theory scaling for entanglement entropy. For open boundary conditions, central charge $c$, and dynamical critical exponent $z$ it reads:
@@ V Conclusions
Overall, we have stressed that addressing finite temperature equilibrium problems requires dedicated algorithms. From a computational perspective, simulating finite-temperature, i.e. mixed many-body quantum states, represents an additional difficulty in comparison to the already computationally demanding zero-temperature quantum many-body simulations. Our algorithm represents an efficient way of obtaining the density matrix and provides adaptability to convert from LPTN geometry, suitable for imaginary time evolution, to TTO geometry, suitable for characterization of the mixed state properties. The procedure is flexible to use beyond Rydberg systems and therefore can be employed as a toolbox for studying finite temperature physics of a general quantum many-body model.

### v2

@@ Abstract
With one or more electrons in a highly excited energy state, Rydberg atoms display strong interparticle interactions Urban et al. (2009); Löw et al. (2012); Weber et al. (2017), making them excellent candidates for programmable quantum devices Wu et al. (2021); Browaeys and Lahaye (2020). With the possibility of arranging individual Rydberg atoms in arrays of arbitrary geometries using optical tweezers Kaufman and Ni (2021); Barredo et al. (2018); Endres et al. (2016); Manetsch et al. (2024); Pause et al. (2024); Schlosser et al. (2023), recent advancements in quantum computing with neutral atoms have demonstrated high-fidelity qubit gates Levine et al. (2019); Zhuo et al. (2022); Evered et al. (2023) and arbitrary connectivity Bluvstein et al. (2022, 2023). Moreover, the progress in analog quantum simulation enables the preparation and control of hundreds of neutral atom qubits, allowing to probe quantum many-body properties directly through measurements Gross and Bloch (2017); Browaeys and Lahaye (2020); Wurtz et al. (2023). Quantum simulators have demonstrated to be successful in exploring the various quantum phases, critical behavior, and underlying universal properties Bernien et al. (2017); Keesling et al. (2019); De Léséleuc et al. (2019); Scholl et al. (2021); Ebadi et al. (2021); Zhang et al. (2024).
@@
The paper is structured as follows: in Section I, we explain the numerical method. Then in Section II, we introduce the Rydberg Hamiltonian that we simulate, and give a short overview of the previous findings regarding the structure of the ground state phase diagram. We show our results in Sections III and IV. Finally, we summarize the conclusions in Section V. Details on the numerical method can be found in Appendices A, B, and C.
@@ I Method
In order to construct an efficient numerical algorithm for studying finite-temperature physics, it is important to choose an appropriate representation of the density matrix. Since the TTO representation is well-suited for the extraction of the otherwise hardly accessible mixed-state properties in a low-temperature regime, here we develop the algorithm for constructing the TTO density matrix for quantum many-body systems at thermal equilibrium. In this section, we first provide a recap of TTOs and their advantages in representing and compressing the density matrix. Then, we provide the definitions of two mixed state entanglement measures and clarify how to extract them from a TTO. Finally, we present the main steps of our numerical algorithm.
@@ I.1 Tree Tensor Operator
where |$\ket{\psi_{j}}$ is the $j^{\mathrm{th}}$ eigenstate of the system’s Hamiltonian with corresponding eigenenergy $E_{j}$, $p_{j}$ are the classical probabilities of finding a system in a state |$\ket{\psi_{j}}$, $Z$ is the partition function, and $\beta=1/T$ with $T$ being temperature (in Boltzmann’s constant units, i.e.,, $k_{B}=1$). Even though, in principle, the number of eigenstates that contribute to the density matrix grows exponentially with the system size, the probabilities $p_{j}$ decay exponentially with the eigenstate’s energy. The temperature factor in the probabilities regulates the flatness of the distribution and one can see that only the low-energy states have a non-negligible contribution to the density matrix in the low-temperature regime (Fig. 1c). In the TTO structure, the middle connecting link corresponds to the sum over eigenstates in Eq. (1), and thus we have direct access to the number of states kept in the density matrix, determined by the corresponding bond dimension $K_{0}$ (Fig. 1a). With the ability to keep only the first $K_{0}$ eigenstates and discard the information about the rest, we keep the computational cost of TTO algorithms low, while retaining the physically accurate approximation. Altogether, we employ two separate bond dimension bounds in the TTO: $m$ as the bond dimension bounding system-to-system correlations and $K_{0}$ as the bound for the mixing bond dimension, i.e., for system-environment correlations.
@@
I.2 Extracting the mixed-state entanglement
As described in Fig. 1b, the complete information about the bipartite entanglement between the system’s halves is contained in the TTO’s root tensor $R$. Thus, $RR^{{\dagger}}$ represents a compressed density matrix from which we can compute the mixed-state entanglement measures.
Here, we compute two measures: the entanglement negativity and the entanglement of formation. The entanglement negativity is defined as
$\begin{split}&E_{N}=\frac{\lVert\hat{\rho}^{\top_{N/2}}\rVert_{1}-1}{2},\end{split}$ (2)
where $\hat{\rho}^{\top_{N/2}}$ is the partial transpose of $\hat{\rho}$ with respect to half of the system and $\lVert A\rVert_{1}$ is the trace norm, $\lVert A\rVert_{1}=Tr(\sqrt{A^{{\dagger}}A})$. The computation of the partial transpose $\hat{\rho}^{\top_{N/2}}$ from a TTO is graphically shown in Fig. 2.
Figure 2: Partial transpose of a TTO density matrix for computing the entanglement negativity. Only the root tensors (purple) are relevant for extracting the entanglement negativity, thus the rest of the TTO can be discarded. Then, the partial transpose with respect to half of the system equals to flipping the corresponding legs (orange and green).
The entanglement of formation is the mixed-state generalization of entanglement entropy. Mathematically, it is defined as a weighted sum of entanglement entropies minimized over all the possible pure-state decompositions of $\hat{\rho}$:
with $\mathcal{S}_{E}$ denoting the entanglement entropy of half-system bipartition. In general, calculating the entanglement of formation is an extremely challenging task, and exact solutions are known only for small system sizes Hill and Wootters (1997); Wootters (1998) or very specific cases Terhal and Vollbrecht (2000); Vollbrecht and Werner (2001). Here, we compute it using the procedure presented in Ref. Arceci et al. (2021), improved as described in Appendix B.
I.3 Obtaining the Tree Tensor Operator density matrix
To access the system’s properties at an arbitrary temperature, we evolve the infinite-temperature density matrix according to the imaginary time evolution Verstraete et al. (2004); Zwolak and Vidal (2004); Weiss et al. (2018). This is possible due to the fact that an infinite-temperature density matrix, $\hat{\rho}_{\infty}$, is proportional to an identity matrix, allowing us to write:
$\begin{split}\hat{\rho}(T)&=e^{-\beta(T)\hat{H}}=e^{-\beta(T)\hat{H}/2}\mathds{1}e^{-\beta(T)\hat{H}/2}\\ &\propto e^{-\beta(T)\hat{H}/2}\hat{\rho}_{\infty}e^{-\beta(T)\hat{H}/2}.\end{split}$ (4)
Figure 3: An algorithm for obtaining the thermal density matrix in the TTO form. An infinite-temperature density matrix is always proportional to an identity matrix, and it is therefore straightforward to write in the LPTN form. We take this infinite-temperature density matrix as an initial state and evolve it via the imaginary time evolution, i.e., consecutively applying the operator $e^{-\mathrm{d}\beta\hat{H}/2}$ on both sides of the LPTN. Here, $\mathrm{d}\beta/2$ is an imaginary time step which, in terms of physical interpretation, represents a small decrease in temperature. This way, we can reach a density matrix at an arbitrarily low temperature. Once we reach the desired temperature, we convert the LPTN into the TTO form with the procedure described in Appendix A.
[…]
@@ I.3 Obtaining the Tree Tensor Operator density matrix
We evolve LPTN in imaginary time using the time-dependent variational principle (TDVP) Haegeman et al. (2011, 2016) algorithm. For details on the TDVP algorithm see Ref. Paeckel et al. (2019); Jaschke et al. (2018). Once we obtain the finite-temperature LPTN density matrix, we convert it to a TTO with an iterative procedure described in Appendix A. The whole procedure for obtaining a TTO thermal density matrix is summarized in Fig. 3.
@@ II Rydberg atom Hamiltonian
To identify the ground state phases in the parameter range of our simulations, we measure the entanglement negativity $E_{N}$ between two halves of the system at temperature $T=0.025\ \Omega$ in Fig. 4. As shown in the analysis in Section III, this temperature is low enough for the system to be approximately in the ground state.
Even though entanglement negativity is constructed as a mixed-state entanglement measure, here we show that it can also be used for quantifying the pure state entanglement, as the phase diagram in Fig. 4 is in line with the existing results obtained via the entanglement entropy Rader and Lauchli (2019); Yu et al. (2022); Maceira et al. (2022). The phase transition lines correspond to maxima in Fig. 4 where one can differentiate three phases in our parameter range: a disordered phase and two crystalline phases. It is known that the transition line between the $\mathds{Z}_{2}$ and disordered phase exhibits the Ising universality behavior. The transition between the $\mathds{Z}_{3}$ and the disordered phase is more complex, and in the parameter range of our simulations, it contains the non-conformal chiral transition line and an isolated point of the three-state Potts universality class.
@@ III Finite temperature phase diagrams
[…]
Recall that the accuracy of our method is limited by the number of excited states we keep in a density matrix, $K_{0}$. This means that, when the state becomes more mixed, the state truncation error becomes larger at fixed $K_{0}$. An easy way of estimating the error induced by discarding the excited states is by tracking the norm loss during the LPTN to TTO conversion. The blue lines in Fig. 5 indicate the temperatures after which the preserved norm falls under a certain factor. As expected, the error increases with the system size for a fixed temperature value. This is because the number of excited states in the spectrum grows exponentially with the system size, and thus, increasing the system size, the number of states $K_{0}$ will represent a smaller ratio of all possible states. A more detailed convergence analysis is carried out in Appendix C.
To get an overview of the temperature resilience of ground state phases over the entire phase diagram, we plot the threshold temperature at which the purity falls below a chosen cutoff value, across the entire Hamiltonian parameter range. The obtained finite-temperature phase diagrams for two different values of threshold purity value, $\gamma_{cut}=\{0.5,0.9\}$, are shown in Fig. 6 and one can see that they outline the features of a ground state phase diagram from Fig. 4. The plot minima correspond to the borders between the phases, and the ground state becomes more stable as we go deeper into the crystalline phases. Again, we observe that the transition from the ground state to a mixture of excited states occurs faster for larger systems, i.e., as we go towards the thermodynamic limit.
@@ IV Mixed-state entanglement at quantum critical points
In this section, we focus on the conformal phase transition points. In particular, we compute the mixed-state entanglement monotones at two points belonging to the Ising universality class and the Potts universality class Yu et al. (2022) (marked with a blue diamond and a green star in Fig. 4). Exploiting the TTO ansatz, we measure the temperature dependence of the bipartite entanglement between the two halves of the system via the entanglement negativity defined in Eq. (2) and the entanglement of formation defined in Eq. (3).
In order to improve the convergence of the results, we truncate the $K_{0}$ dimension even further and keep the first $K_{0}=10$ states in the density matrix in the computation of entanglement of formation. Note that, as discussed in Section III, reducing the number of states in the density matrix increases the error in the high-temperature region. Therefore, we carry out the entanglement of formation analysis only in the low-temperature region in which the norm loss is smaller than $1\%$.
@@
The next step is to analyze and quantify finite-size effects, i.e., how entanglement scales with system size and temperature. The numerical results on the entanglement of formation $E_{F}$ in low-temperature Ising and Luttinger liquid XXZ model Arceci et al. (2021) predicted that in the vicinity of the conformal quantum critical point, $E_{F}$ should follow the logarithmic scaling. This scaling relation is the finite-temperature extension of the conformal field theory scaling for entanglement entropy. For open boundary conditions, central charge $c$, and dynamical critical exponent $z$ it reads:
@@ V Conclusions
Overall, we have stressed that addressing finite temperature equilibrium problems requires dedicated algorithms. From a computational perspective, simulating finite-temperature, i.e., mixed many-body quantum states, represents an additional difficulty in comparison to the already computationally demanding zero-temperature quantum many-body simulations. Our algorithm represents an efficient way of obtaining the density matrix and provides adaptability to convert from LPTN geometry, suitable for imaginary time evolution, to TTO geometry, suitable for characterization of the mixed state properties. The procedure is flexible to use beyond Rydberg systems and therefore can be employed as a toolbox for studying finite temperature physics of a general quantum many-body model. We expect that the most efficient approach to extending the framework towards two-dimensional system analysis is by mapping a model to a corresponding one-dimensional one Cataldi et al. (2021). Another alternative is the conversion of a projected entangled-pair operator Pirvu et al. (2010); Alhambra and Cirac (2021), i.e., a density matrix in projected entangled-pair state form, into a TTO.
@@ Appendix B Entanglement of formation minimization
Appendix C Convergence analysis
In this section, we show how the convergence of the LPTN to TTO conversion algorithm depends on the bond dimension $m$. In the analysis, we study the system of $N=31$ sites at the Ising and the Potts critical points. We choose to showcase the largest system size and critical states because these points are the computationally most demanding task of this work. We generate LPTN states with imaginary time evolution timestep $dt=0.025$, in the temperature regime $T/\Omega=[0.02,0.4]$. To provide a meaningful analysis of the convergence of the LPTN to TTO conversion scheme, we make sure that the LPTN state from which we start the conversion is converged in the bond dimension. By increasing the bond dimension from $m=30$ to $m=60$, values of LPTN energy remain unchanged up to 8th digit for every temperature point. Therefore, we assume that the states of these bond dimensions are converged.
In the LPTN to TTO conversion, we fix the bond dimension $K_{0}=100$, i.e., we keep the first $100$ states in a density matrix. We compare the data for bond dimensions $m=[30$, $40$, $50$, $60$]. To measure the accuracy of the conversion, we track the state’s norm loss due to singular value truncations during the conversion (Fig. 9a) and compare the energy of the resulting TTO with the corresponding energy of the LPTN (Fig. 9b). To give an intuition on how this reflects on the convergence of entanglement results, we show the convergence of entanglement negativity for different bond dimensions $m$ (Fig. 9c).
Before continuing to the convergence analysis results, it is useful to highlight the different roles of the bond dimensions $m$ and $K_{0}$ in a TTO. Interpretation of $K_{0}$ is straightforward, as $K_{0}$ is simply the final number of pure states we keep in the density matrix. Interpretation of $m$ is, however, much more subtle: if one thinks of a TTO as multiple pure TTN states stacked together, TTO’s $m$ contains entanglement distributed over many pure states. Thus, a required $m$ depends on the amount of entanglement of each pure state and overall mixedness of a state - more entanglement in each pure state requires larger $m$, but also larger mixedness requires larger $m$ as there are more pure states to be covered. The distribution between the two contributions within $m$ is not transparent.
Back to convergence, as already discussed in Sec. III, increasing the temperature means increasing the number of excited states that contribute to the density matrix. Therefore, the accuracy of results decreases as the number of states $K_{0}$ becomes insufficient to describe a physical state. This reflects directly on the norm loss in Fig. 9a. As we can see, the norm loss curves do not significantly vary with the bond dimension $m$, thus for these bond dimensions it is the choice of $K_{0}$ that sets the limit to the overall accuracy of the results. Note that the norm loss would be influenced by $m$ for smaller values of $m$. We denote the regime in which the norm loss is smaller than 1% with a black vertical line, physically representing an approximate temperature at which $K_{0}=100$ states are enough to describe a thermal density matrix. Therefore, we expect our results to be accurate below this threshold temperature. Since the norm loss and the threshold temperature depend on the system size and the point in the phase diagram, we show the norm loss lines as well in Fig. 5.
By looking at the energy curves in Fig. 9b, the TTO results in the region below the threshold temperature coincide with the LPTN reference values. Also, the entanglement negativity in Fig. 9c is well-converged with $m$ within this region. The energy and the negativity at the Potts point and the negativity at the Ising point appear not to be converged with $m$ in the temperature regions above the threshold temperature. We attribute this to the above-discussed effect where larger mixedness of the system requires a larger bond dimension. Nevertheless, this region is not of interest because the accuracy is limited by $K_{0}$ and not $m$, therefore we do not expect accurate results there. An additional remark regarding the convergence results is that, even though the $m=30$ energy curve for the Potts point coincides with the reference LPTN curve in the plot, we expect this to be an accidental artifact of unconvergence in $m$.
Figure 9: Convergence of LPTN to TTO conversion algorithm with the bond dimension $m$. Norm loss during the conversion (a), energy (b), and entanglement negativity (c) are computed at the Ising and the Potts critical points of the Rydberg model for $N=31$ sites. The bond dimension $m$ of the TTO and the corresponding LPTN state are the same. The bond dimension $K_{0}$ is fixed to $K_{0}=100$. The vertical line in all plots denotes a threshold temperature below which the norm loss during the LPTN to TTO conversion is smaller than 1%. The LPTN energy curve (blue curve) in (b) represents the reference values computed for $m=60$.
In this analysis, we keep the bond dimension $m$ the same for both LPTN and TTO states, but in general, they do not have to be the same. A good choice is that $m_{\mathrm{TTO}}$ is larger or equal to $m_{\mathrm{LPTN}}$. Overall, we suggest the following strategy for picking the bond dimensions for LPTN to TTO conversion of thermal states:
  1. 1.
    Choose $m_{\mathrm{LPTN}}$ such that LPTN states are converged.
  2. 2.
    Pick $K_{0}$ and tune $m_{\mathrm{TTO}}$ such that the norm loss has converged in $m_{\mathrm{TTO}}$.
  3. 3.
    Read out the threshold temperature below which the norm loss is smaller than the desired precision.
  4. 4.
    If the threshold temperature is too low for the purpose of the simulation, repeat the steps $1.-3$ with a larger $K_{0}$.
  5. 5.
    Double-check if the properties have converged with $m_{\mathrm{TTO}}$ in region below the threshold temperature.
Note that, in practice, we cannot tune the threshold temperature to drastically higher, as the required number of states $K_{0}$ soon becomes too large due to energies in the spectrum being dense or clustered in bands. In general, the available temperature range is of the order of magnitude of the gap between the ground and the first excited state.